# The Statistical Fairness-Accuracy Frontier

### v1

@@ 2.1 Fairness-Accuracy Frontier
We next adopt the definition of the fairness-accuracy (FA) Pareto frontier from Liang et al. [2021]. To do so, we first revisit the definition of FA dominance.
@@
There are three key points in understanding the fairness–accuracy (FA) frontier: the two points corresponding to the prediction functions that achieve the best accuracy for each group, and the point that achieves the best fairness guarantee. More formally, for any $g\in\mathcal{G}$, let $f_{g}$ be defined as
@@
It is straightforward to verify that the pairs of risks corresponding to all three functions lie on the FA frontier. In fact, the endpoints of the FA frontier are two of the three points associated with $\{f_{r},f_{b},f_{f}\}$. Liang et al. [2021] define a distribution $P$ as group-skewed if the point corresponding to $f_{f}$ lies at one of the endpoints of the frontier. In contrast, $P$ is said to be group-balanced if the endpoints correspond to $f_{r}$ and $f_{b}$, and the most fair point $f_{f}$ lies between them. An illustration of group-skewed and group-balanced distributions is provided in Figure 3.
@@ 2.2 Empirical Fairness-Accuracy Frontier
In practice, the true distribution $P$ is unknown, so $f_{\lambda}$ is not computationally tractable. Instead, we may learn an empirical estimate $\widehat{f}_{\lambda}\in\mathcal{H}$ based on the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. The error of this estimation, known as the excess risk, is given by $\mathcal{R}_{\lambda}(\widehat{f}_{\lambda})-\mathcal{R}_{\lambda}(f_{\lambda})$, and is nonnegative, since $f_{\lambda}$ minimizes $\mathcal{R}_{\lambda}(\cdot)$.
Note that the corresponding pair of risks $(\mathcal{R}_{r}(\widehat{f}_{\lambda}),\mathcal{R}_{b}(\widehat{f}_{\lambda}))$ will always lie in $\mathcal{E}(\mathcal{H})$, and may fall strictly in its interior. This implies that the empirical fairness-accuracy frontier contracts inward relative to the FA frontier $\mathcal{F}(\mathcal{H})$. That said, for any estimator $\widehat{f}_{\lambda}$, the corresponding excess risk quantifies how much the frontier is pushed inward in the direction orthogonal to the line $\lambda\mathcal{R}_{r}+(1-\lambda)\mathcal{R}_{b}=\text{constant}$. An illustration of this phenomenon is shown in Figure 4(b).
Now, assuming that the distribution $P$ belongs to a class of distributions $\mathcal{P}$, we can take a worst-case approach by finding an estimator $\widehat{f}_{\lambda}$ that minimizes the worst-case excess risk. More formally, we aim to solve the following problem:
@@
where the expectation is taken with respect to the draw of the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. In the remainder of the paper, we focus on the linear regression problem. We present a lower bound for the worst-case excess risk in (4), and develop an estimator that matches this lower bound with respect to many of the problem’s parameters. We also discuss how the risks for the two groups change when this optimal estimator is used.
@@ 3 Linear Model Class
To provide theoretical bounds on the empirical FA frontier, we restrict attention to the class of linear models:
@@
A canonical example of a distribution in $\mathcal{P}_{\text{linear}}(\sigma^{2})$ is the well-specified linear model with homoskedastic Gaussian noise $\varepsilon_{g}\sim\mathcal{N}(0,\sigma^{2})$, independent of $X$.
In what follows, we take $\mathcal{H}$ as the class of linear models, i.e., $\{f_{\beta}(x)=x^{\top}\beta:\beta\in\mathbb{R}^{d}\}$. In addition, we take the squared loss function $\ell(f_{\beta}(X),Y)=(\beta^{\top}X-Y)^{2}$, and also simplify the notation of the risks to $\mathcal{R}_{g}(f_{\beta})$ and $\mathcal{R}_{\lambda}(f_{\beta})$ to $\mathcal{R}_{g}(\beta)$ and $\mathcal{R}_{\lambda}(\beta)$, respectively. We write $\mathbb{E}_{g}$ and $\mathbb{P}_{g}$ to denote expectation and probability, respectively, under the conditional distribution $P(X,Y|G=g)$.
@@ 3.1 Assumptions
The above assumption requires that $n_{g}\geq d$. Given this condition, and as highlighted by Mourtada [2022, Fact 1], 2 is equivalent to either of the following two properties: (i) the distribution $P(X\mid G=g)$ does not assign positive probability to any linear hyperplane; and (ii) the ordinary least squares (OLS) estimator for each group, defined as
@@
This condition, adopted from prior work (e.g., Koltchinskii and Mendelson [2015], Mourtada [2022]), provides lower-tail control of the sample covariance spectrum, ensuring that the empirical covariance does not degenerate in directions of low variance of the covariates. This assumption holds for multivariate Gaussian distributions with $\alpha_{g}=1$, since in that case $\theta^{\top}X$ is Gaussian with variance $\|\theta\|_{\Sigma_{g}}^{2}$. Furthermore, Rudelson and Vershynin [2015] show that this condition holds with $\alpha_{g}=1$ for covariates with independent coordinates and bounded densities.
@@
$\left\lVert X^{\top}u\right\rVert_{\psi_{2}}\leq K_{g}~\big{\|}u\big{\|}_{\Sigma_{g}},$
where the $\psi_{2}$-norm is defined conditional on the group identity being $g$.
@@ 3.2 Empirical Estimators
Let $\beta_{\lambda}$ denote the optimal predictor corresponding to the weighted loss $\mathcal{R}_{\lambda}(\beta)$, i.e.,
@@
where $\beta$ is a function of the datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. [^2] Moving forward, we omit the dependence of the expectation on the datasets when it is clear from the context. For the cases $\lambda=0$ and $\lambda=1$, which reduce to the single-group setting (say, group $g$), Mourtada [2022] shows that the OLS estimator $\widehat{\beta}_{g}$, defined in (6), is the optimal worst-case estimator with respect to the minimax criterion in (10), when $\mathcal{P}$ is taken to be the linear model class $\mathcal{P}_{\text{linear}}(\sigma^{2})$. Interestingly, he also shows that knowledge of the covariate distribution $P_{X}^{g}$ does not affect this optimality: the OLS estimator remains minimax optimal even when the supremum is taken over the more restricted class $\mathcal{P}_{\text{linear}}(P_{X}^{g},\sigma^{2})$.
This, naturally, leads us to consider the empirical version of $\mathcal{R}_{\lambda}(\cdot)$, denoted by $\widehat{\mathcal{R}}_{\lambda}(\beta)$, which is given by
@@ 4 Estimation with Known Covariances
with $\Sigma_{\lambda}=\lambda\Sigma_{r}+(1-\lambda)\Sigma_{b}$, and $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}$, as defined in (8) and (6), respectively. This estimator maintains the structure of the optimal predictor $\beta_{\lambda}$ but replaces the true cross moment $\nu_{g}=\Sigma_{g}\beta_{g}$ with the empirical quantity $\Sigma_{g}\widehat{\beta}_{g}$ for each group $g$.
@@ 4.1 The Optimal Estimator
As Proposition 1 shows, the excess risk of an estimator can be decomposed into the sum of its distances from the optimal predictor $\beta_{\lambda}$ under the Mahalanobis norms induced by $\Sigma_{r}$ and $\Sigma_{b}$. The following result shows that, for each group $g$, the worst-case distance $\|\beta-\beta_{\lambda}\|_{g}^{2}$ is minimized by $\tilde{\beta}_{\lambda}$, thereby implying the minimax optimality of this estimator.
@@ 4.2 Bounding the Excess Risk
The constant $C’_{g}$ arises from the bound on the minimum eigenvalue of the sample covariance, which is used here to bound the trace of $\widehat{\Sigma}_{r}^{-1}$ and $\widehat{\Sigma}_{b}^{-1}$, as given in Mourtada [2022, Theorem 4]. However, as highlighted in Mourtada [2022, Remark 6]—based on results from Wu and Verdú [2012] and Edelman [1988]—this constant can be significantly reduced for Gaussian distributions. In particular, if the covariate distributions are Gaussian, $C’_{g}$ can be replaced by a constant whose limit, when ${d}/{n_{g}}\to h\in(0,1)$, is upper bounded by
@@
Per-group estimation error: finite-sample effects can shift fairness–accuracy trade-offs
@@
It is worth noting that the results above are in expectation over the draw of datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$, and not for a particular realization. For the excess risk with respect to the weighted loss, i.e., $\mathcal{R}_{\lambda}(\tilde{\beta}_{\lambda})-\mathcal{R}_{\lambda}(\beta_{\lambda})$, however, each realization is also positive. This means that the empirical estimator always increases the loss function $\mathcal{R}_{\lambda}(\cdot)$, simply because $\beta_{\lambda}$ is its minimizer. By contrast, this monotonicity does not hold for per-group risks: one group’s estimation error $\mathcal{R}_{g}(\cdot)$ might decrease under the empirical estimator relative to $\beta_{\lambda}$, even though the other group’s risk increases enough to make the weighted objective worse. The following result formalizes this distinction.
@@
The proof is given in Section A.3.3. In summary, this result follows from the decomposition established in Proposition 1:
@@ 5 Estimation with Unknown Covariance
We next consider the setting in which the distributions of the covariates are unknown and must be estimated from data. Here, our proposed estimator is the OLS estimator (12), which, as we recall, is given by
$\widehat{\beta}_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}=(\lambda\widehat{\Sigma}_{r}+(1-\lambda)\widehat{\Sigma}_{b})^{-1}(\lambda\widehat{\nu}_{r}+(1-\lambda)\widehat{\nu}_{b}).$
@@ 5.1 The Bias-Variance Decomposition
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]\leq\mathcal{V}_{g}(\lambda)+\mathcal{B}_{g}(\lambda),$ (21)
@@ 5.2 Upper and Lower Bounds for the Bias and Variance Terms
where $\mathcal{V}_{g}(\lambda)$ is the variance term, as defined in Proposition 3, and $C^{\prime}_{g}$ is given in Corollary 2 and Remark 3. Moreover, assuming $n_{g}\geq\sigma^{2}/(B\rho_{g}^{2})$ for $g\in\mathcal{G}$, we have
@@
The choice of the perturbations $\{u_{i}\}$ is key here. Since the bias term has the form $M(\beta_{r}-\beta_{b})$ for some matrix $M$, its magnitude is maximized when $\beta_{r}-\beta_{b}$ is closely aligned with an eigenvector of $M$ corresponding to its largest eigenvalue. The construction of $u_{i}$ ensures that each inner product has a fixed nontrivial inner product with both $\beta_{r}-\beta_{b}$ and the canonical basis vector $e_{i}$. This alignment guarantees that we perturb the covariance matrix along all coordinate directions while maintaining proximity to $\beta_{r}-\beta_{b}$.
@@
We can pose the same question as in the previous section: given a fixed budget on $n_{r}+n_{b}$, what is the optimal way to choose these two parameters to minimize the excess risk? Theorem 2 suggests that, for minimizing the variance term, the answer is similar to the case of known covariances. In fact, similar to (17), we should choose ${n_{r}}/{n_{b}}=(\lambda\rho_{r})/((1-\lambda)\rho_{b})$ to minimize the variance term. In contrast, Theorem 3 shows that minimizing the bias term requires a balanced design, namely, $n_{r}=n_{b}$.
@@ 5.3 Per-Group Estimation Errors
However, the bound is in fact not symmetric. As stated after the proposition, it changes from $\lambda(1-\lambda)^{2}$ for group $r$ to $\lambda^{2}(1-\lambda)$ for group $b$. This bias effect is larger for group $r$ when $\lambda$ is close to 0 and for group $b$ when $\lambda$ is close to $1$, meaning that the second term increases in absolute value for one group when the other group is prioritized in the overall loss function.

### v2

@@ 2.1 Fairness-Accuracy Frontier
We next adopt the definition of the fairness-accuracy (FA) Pareto frontier from liang2021algorithm. To do so, we first revisit the definition of FA dominance.
@@
Insight into the shape of the fairness-accuracy (FA) frontier can be obtained by focusing on three key points: the two points corresponding to the prediction functions that achieve the best accuracy for each group, and the point that achieves the best fairness guarantee. More formally, for any $g\in\mathcal{G}$, let $f_{g}$ be defined as
@@
It is straightforward to verify that the pairs of risks corresponding to all three functions lie on the FA frontier. In fact, the endpoints of the FA frontier are two of the three points associated with $\{f_{r},f_{b},f_{f}\}$. liang2021algorithm define a distribution $P$ as group-skewed if the point corresponding to $f_{f}$ lies at one of the endpoints of the frontier. In contrast, $P$ is said to be group-balanced if the endpoints correspond to $f_{r}$ and $f_{b}$, and the most fair point $f_{f}$ lies between them. An illustration of group-skewed and group-balanced distributions is provided in Figure 3.
@@ 2.2 Empirical Fairness-Accuracy Frontier
In practice, the true distribution $P$ is unknown, so $f_{\lambda}$ is not available. Instead, we begin by obtaining an empirical estimate $\widehat{f}_{\lambda}\in\mathcal{H}$ based on the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. The error associated with this estimate, known as the excess risk, is given by $\mathcal{R}_{\lambda}(\widehat{f}_{\lambda})-\mathcal{R}_{\lambda}(f_{\lambda})$. It is nonnegative, since $f_{\lambda}$ minimizes $\mathcal{R}_{\lambda}(\cdot)$.
Note that the corresponding pair of risks $(\mathcal{R}_{r}(\widehat{f}_{\lambda}),\mathcal{R}_{b}(\widehat{f}_{\lambda}))$ will always lie in $\mathcal{E}(\mathcal{H})$, and may fall strictly in its interior. Moreover, for any estimator $\widehat{f}_{\lambda}$, the corresponding excess risk quantifies how much the frontier is pushed inward in the direction orthogonal to the line $\lambda\mathcal{R}_{r}+(1-\lambda)\mathcal{R}_{b}=\text{constant}$. An illustration of this phenomenon is shown in Figure 4(b).
Now, assuming that the distribution $P$ belongs to a class of distributions $\mathcal{P}$, we can proceed without knowledge of $P$ by finding an estimator $\widehat{f}_{\lambda}$ that minimizes the worst-case excess risk. More formally, we aim to solve the following problem:
@@
where the expectation is taken with respect to the draw of the dataset $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. In the remainder of the paper, we study this minimax problem in the setting of linear regression. We present a lower bound for the worst-case excess risk in (4), and develop an estimator that matches this lower bound with respect to many of the problem’s parameters. We also discuss how the risks for the two groups change when this optimal estimator is used.
@@ 3 Linear Model Class
To study theoretical bounds on the empirical FA frontier, we restrict attention to the class of linear models:
@@
A canonical example of a distribution in $\mathcal{P}_{\text{linear}}(\sigma^{2})$ is the well-specified linear model with homoskedastic Gaussian noise, $\varepsilon_{g}\sim\mathcal{N}(0,\sigma^{2})$, independent of $X$.
In what follows, we take $\mathcal{H}$ as the class of linear models, i.e., $\{f_{\beta}(x)=x^{\top}\beta:\beta\in\mathbb{R}^{d}\}$. In addition, we consider the squared loss function $\ell(f_{\beta}(X),Y)=(\beta^{\top}X-Y)^{2}$, and also simplify the notation of the risks $\mathcal{R}_{g}(f_{\beta})$ and $\mathcal{R}_{\lambda}(f_{\beta})$ to $\mathcal{R}_{g}(\beta)$ and $\mathcal{R}_{\lambda}(\beta)$, respectively. We write $\mathbb{E}_{g}$ and $\mathbb{P}_{g}$ to denote expectation and probability, respectively, under the conditional distribution $P(X,Y|G=g)$.
@@ 3.1 Assumptions
This assumption requires that $n_{g}\geq d$. Given this condition, and as highlighted by mourtada2022exact, 2 is equivalent to either of the following two properties: (i) the distribution $P(X\mid G=g)$ does not assign positive probability to any linear hyperplane; and (ii) the ordinary least squares (OLS) estimator for each group, defined as
@@
This condition, adopted from prior work (e.g., koltchinskii2015bounding, mourtada2022exact), provides lower-tail control for the sample covariance spectrum, ensuring that the empirical covariance does not degenerate in directions of low variance of the covariates. This assumption holds for multivariate Gaussian distributions with $\alpha_{g}=1$, since in that case $\theta^{\top}X$ is Gaussian with variance $\|\theta\|_{\Sigma_{g}}^{2}$. Furthermore, rudelson2015small show that this condition holds with $\alpha_{g}=1$ for covariates with independent coordinates and bounded density.
@@
$\mathbb{P}_{g}\left(\left|X^{\top}u\right|\geq t\left\lVert u\right\rVert_{\Sigma_{g}}\right)\leq 2\exp\left(\frac{-t^{2}}{K_{g}^{2}}\right).$
@@ 3.2 Empirical Estimators
Let $\beta_{\lambda}$ denote the optimal predictor corresponding to the weighted loss $\mathcal{R}_{\lambda}(\beta)$:
@@
where $\beta$ is a function of the datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$. [^2] Moving forward, we omit the dependence of the expectation on the datasets when it is clear from the context. For the cases $\lambda=0$ and $\lambda=1$, which reduce to the single-group setting (say, group $g$), mourtada2022exact shows that the OLS estimator $\widehat{\beta}_{g}$, defined in (6), is the optimal worst-case estimator with respect to the minimax criterion in (10), when $\mathcal{P}$ is taken to be the linear model class $\mathcal{P}_{\text{linear}}(\sigma^{2})$. Interestingly, he also shows that knowledge of the covariate distribution $P_{X}^{g}$ does not affect this optimality: the OLS estimator remains minimax optimal even when the supremum is taken over the more restricted class $\mathcal{P}_{\text{linear}}(P_{X}^{g},\sigma^{2})$.
This naturally leads us to consider the empirical version of $\mathcal{R}_{\lambda}(\cdot)$, denoted by $\widehat{\mathcal{R}}_{\lambda}(\beta)$, which is given by
@@ 4 Estimation with Known Covariances
with $\Sigma_{\lambda}=\lambda\Sigma_{r}+(1-\lambda)\Sigma_{b}$, and $\widehat{\beta}_{g}=\widehat{\Sigma}_{g}^{-1}\widehat{\nu}_{g}$, as defined in (8) and (6), respectively. This estimator maintains the structure of the optimal predictor $\beta_{\lambda}$ but replaces the true cross-moment $\nu_{g}=\Sigma_{g}\beta_{g}$ with the empirical quantity $\Sigma_{g}\widehat{\beta}_{g}$ for each group $g$.
@@ 4.1 The Optimal Estimator
As Proposition 1 shows, the excess risk of an estimator can be decomposed into the sum of its distances from the optimal predictor $\beta_{\lambda}$ under the Mahalanobis norms induced by $\Sigma_{r}$ and $\Sigma_{b}$. The following result shows that, for each group $g$, the worst-case distance $\|\beta-\beta_{\lambda}\|_{\Sigma_{g}}^{2}$ is minimized by $\tilde{\beta}_{\lambda}$, thereby implying the minimax optimality of this estimator.
@@ 4.2 Bounding the excess risk
The constant $C’_{g}$ arises from the bound on the minimum eigenvalue of the sample covariance, which is used here to bound the trace of $\widehat{\Sigma}_{r}^{-1}$ and $\widehat{\Sigma}_{b}^{-1}$, as given in mourtada2022exact. However, as highlighted in mourtada2022exact—based on results from wu2012optimal and edelman1988eigenvalues—this constant can be significantly reduced for Gaussian distributions. In particular, if the covariate distributions are Gaussian, $C’_{g}$ can be replaced by a constant whose limit, when ${d}/{n_{g}}\to h\in(0,1)$, is upper bounded by
@@
Per-group estimation error—finite-sample effects can shift fairness–accuracy trade-offs:
@@
It is worth noting that the results above are computed in expectation over the random draw of datasets $\{\mathcal{S}_{r},\mathcal{S}_{b}\}$, and are not for a particular realization. For the excess risk with respect to the weighted loss, i.e., $\mathcal{R}_{\lambda}(\tilde{\beta}_{\lambda})-\mathcal{R}_{\lambda}(\beta_{\lambda})$, however, each realization is also positive. This means that the empirical estimator always increases the loss function $\mathcal{R}_{\lambda}(\cdot)$, simply because $\beta_{\lambda}$ is its minimizer. By contrast, this monotonicity does not hold for per-group risks: one group’s estimation error $\mathcal{R}_{g}(\cdot)$ might decrease under the empirical estimator relative to $\beta_{\lambda}$, even though the other group’s risk increases enough to make the weighted objective worse. The following result formalizes this distinction.
@@
The proof is given in Section A.3.3. This result follows from the decomposition established in Proposition 1:
@@ 5 Estimation with Unknown Covariance
We next consider the setting in which the distributions of the covariates are unknown and must be estimated from data. Here, our proposed estimator is the OLS estimator (12), which, as we recall, is given by:
$\widehat{\beta}_{\lambda}=\widehat{\Sigma}_{\lambda}^{-1}\widehat{\nu}_{\lambda}=(\lambda\widehat{\Sigma}_{r}+(1-\lambda)\widehat{\Sigma}_{b})^{-1}(\lambda\widehat{\nu}_{r}+(1-\lambda)\widehat{\nu}_{b}).$ (21)
@@ 5.1 The Bias-Variance Decomposition
$\mathbb{E}\left[\left\lVert\widehat{\beta}_{\lambda}-\beta_{\lambda}\right\rVert_{\Sigma_{g}}^{2}\right]=\mathcal{V}_{g}(\lambda)+\mathcal{B}_{g}(\lambda),$ (22)
@@ 5.2 Upper and lower bounds for the bias and variance terms
where $\mathcal{V}_{g}(\lambda)$ is the variance term, as defined in Proposition 3, and $C^{\prime}_{g}$ is given in Corollary 2 and Remark 3. Moreover, assuming $n_{g}\geq\sigma^{2}d/(B\rho_{g}^{2})$ for $g\in\mathcal{G}$, we have
@@
The choice of the perturbations $\{u_{i}\}$ is key here. Since the bias term has the form $M(\beta_{r}-\beta_{b})$ for some matrix $M$, its magnitude is maximized when $\beta_{r}-\beta_{b}$ is closely aligned with an eigenvector of $M$ corresponding to its largest eigenvalue. Hence we choose the perturbations so that this alignment holds, i.e., the construction of $u_{i}$ ensures that it has a positive and constant inner product with both $\beta_{r}-\beta_{b}$ and the canonical basis vector $e_{i}$. This alignment guarantees that we perturb the covariance matrix along all coordinate directions while maintaining proximity to $\beta_{r}-\beta_{b}$.
@@
We can pose the same question as in the previous section: given a fixed budget on $n_{r}+n_{b}$, what is the optimal way to choose these two parameters to minimize the excess risk? Theorem 2 suggests that, for minimizing the variance term, the answer is similar to the case of known covariances. In fact, similar to (17), we should choose ${n_{r}}/{n_{b}}=(\lambda\rho_{r})/((1-\lambda)\rho_{b})$ to minimize the variance term. In contrast, Theorem 3 shows that minimizing the bias term requires a balanced design, namely, $n_{r}=n_{b}$ (under the assumption of $K_{r}=K_{b}$).
@@ 5.3 Per-Group estimation errors
However, the bound is in fact not symmetric. As stated after the proposition, it changes from $\lambda(1-\lambda)^{2}$ for group $r$ to $\lambda^{2}(1-\lambda)$ for group $b$. This bias effect is larger for group $r$ when $\lambda$ is close to zero and for group $b$ when $\lambda$ is close to one, meaning that, at those endpoints, the term is larger in absolute value for the group that is not prioritized.
@@
6 Uniform Frontier Bounds
So far, our analysis has provided finite-sample guarantees for a single fixed fairness–accuracy preference $\lambda$. In practice, however, a designer may wish to consider many values of $\lambda$ in order to select a data-driven choice $\widehat{\lambda}$. Because the empirical errors for different $\lambda$ values are computed using the same datasets, these errors are highly correlated. One approach is to use a union bound, but even then the designer can only afford to check finitely many values of $\lambda$, since the error would accumulate and requesting additional $\lambda$ values would become “costly.”
In this section, we show how to circumvent this issue by obtaining a uniform guarantee over the entire empirical frontier. We achieve this by viewing the empirical error of the FA frontier as a stochastic process indexed by $\lambda\in[0,1]$. This result follows from the construction of a Doob martingale to show a bounded-difference property of the process followed by a chaining argument.
Defining the population and empirical error pairs across groups,
$\mathcal{R}_{\lambda}\coloneq\left(\mathcal{R}_{r}(\beta_{\lambda}),\mathcal{R}_{b}(\beta_{\lambda})\right),\qquad\widehat{\mathcal{R}}_{\lambda}\coloneq\left(\widehat{\mathcal{R}}_{r}(\widehat{\beta}_{\lambda}),\widehat{\mathcal{R}}_{b}(\widehat{\beta}_{\lambda})\right),$
we define the frontier error as a stochastic process:
$\Delta_{\lambda}\coloneq\widehat{\mathcal{R}}_{\lambda}-\mathcal{R}_{\lambda}\in\mathbb{R}^{2},\quad\lambda\in[0,1].$
We will assume that the noise in the linear model given by Definition 3 is subgaussian.
Assumption 5 (Subgaussian noise).
For both groups $g\in\{r,b\}$, the noise $\varepsilon$ is conditionally mean-zero and $\sigma_{\varepsilon}$-subgaussian, independent of $X$.
The following theorem establishes a high-probability uniform bound on the deviation of the empirical fairness-accuracy frontier from its population counterpart across all possible fairness-accuracy preferences $\lambda\in[0,1]$ for the unknown-covariance estimator $\widehat{\beta}$ defined in (21).
Theorem 4.
Suppose Assumptions 1-5 hold. Then, for any fixed $\alpha>2$, there exists a constant $C_{u}$, such that, for every $\delta\in(0,1)$, with probability at least $1-\delta-\eta$, we have
$\sup_{\lambda\in[0,1]}\left\lVert\Delta_{\lambda}\right\rVert_{2}\leq C_{u}\log(\frac{1}{\delta})\left(\frac{\log(n_{r})^{5}}{\sqrt{n_{r}}}+\frac{\log(n_{b})^{5}}{\sqrt{n_{b}}}\right),$
with
$\eta=(n_{r}+n_{b})^{-C\alpha}+\left(C_{r}^{\prime}/2\right)^{\alpha_{r}n_{r}/6}+\left(C_{b}^{\prime}/2\right)^{\alpha_{b}n_{b}/6}.$
The constant $C_{u}$ depends on the problem’s parameters, and the exact dependence is provided in the appendix.
Theorem 4 provides control over the empirical error process with high probability uniformly over the entire frontier, that is, for all $\lambda\in[0,1]$. This result demonstrates that a designer implementing the frontier in finite samples, will, with high probability, land within a confidence band around the population frontier. In particular, with probability at least $1-\delta-\eta$, the empirical curve lies within a distance that scales as $1/\sqrt{n_{r}}+1/\sqrt{n_{b}}$ with the sample sizes of both groups (ignoring logarithmic terms).
We sketch the proof of the uniform empirical frontier bound in Theorem 4 here; the full proof appears in Section A.5. First, we show that, with high probability, features and outcomes are bounded and empirical covariances are well-conditioned, ensuring stability of the estimator. Next, we show that $\widehat{\beta}_{\lambda}$ is Lipschitz in $\lambda$ which implies Lipschitzness in the empirical risks. We then analyze single sample perturbations (where we replace one observation with an i.i.d draw) to obtain a bounded-difference property. We use a Doob martingale argument with Hoeffding’s lemma to establish subgaussianity of the increments in $|\lambda-\lambda^{\prime}|$ for the scalarized process. Chaining via Dudley’s integral inequality converts the increment control to a uniform bound over all $\lambda$, and an $\varepsilon$-net argument over all unit vectors brings the argument back to the two-dimensional empirical error process across the frontier. Finally, we de-condition using the high-probability truncation event to yield the result.
Since the bound given by Theorem 4 holds simultaneously for all $\lambda\in[0,1]$, a designer may select $\widehat{\lambda}$ adaptively after inspecting the empirical frontier while enjoying the same high-probability guarantee. Moreover, the result strengthens the guarantees provided under a naïve union bound argument on a finite grid: the bound holds for the entire interval $[0,1]$ simultaneously, including values not explicitly evaluated, so the designer may refine their search at no additional cost.
7 Experiments
In this section, we conduct synthetic experiments to demonstrate how well the estimators $\tilde{\beta}_{\lambda}$ and $\widehat{\beta}_{\lambda}$, for the known and unknown covariance settings, respectively, perform relative to the population quantity across the FA frontier.
We construct a two-group Gaussian linear model in $d$ dimensions with isotropic covariances $\Sigma_{g}=\rho_{g}^{2}I_{d}$. For each group $g\in\{r,b\}$, features $X_{g}\sim\mathcal{N}(0,\Sigma_{g})$, and outcomes $Y_{g}=X_{g}^{\top}\beta_{g}+\varepsilon$ with $\varepsilon\sim\mathcal{N}(0,\sigma^{2})$. The pair $(\beta_{r},\beta_{b})$ is constructed via a dense random perturbation scaled so that the heterogeneity $\left\lVert\beta_{b}-\beta_{r}\right\rVert_{2}$ is fixed across iterations. We evaluate (i) the deterministic population frontier from $\beta_{\lambda}$ using the true $\Sigma_{g}$ and $\beta_{g}$, (ii) the empirical frontier in the known $\Sigma_{g}$ case from $\tilde{\beta}_{\lambda}$ given in (13), and (iii) the empirical frontier in the unknown $\Sigma_{g}$ case from $\widehat{\beta}_{\lambda}$ given in (21).
[…]
Figure 5: We sweep $\lambda$ on a uniform grid of 5050 points in $[0,1]$. For each $\lambda$, we run 100 Monte Carlo repetitions; in each iteration, we draw data $\mathcal{S}_{g}=\{(X_{g},Y_{g})\}_{i=1}^{n_{g}}$ from the group-$g$ Gaussian linear model with $\sigma^{2}=1$, build the estimators $\tilde{\beta}_{\lambda},\widehat{\beta}_{\lambda}$, and calculate the corresponding risk pair $(\mathcal{R}_{r},\mathcal{R}_{b})$ for each estimator. We plot the mean empirical frontiers across repetitions and depict the empirical contraction for the choice of $\lambda=0.5$ for two regimes: (a) Equal samples: $n_{r}=n_{b}=30,d=12,\rho_{r}=1,\rho_{b}=2.5$, (b) Equal covariance: $n_{r}=50,n_{b}=30,d=12,\rho_{r}=\rho_{b}=1.75$.
Figure 5 isolates the effect of heterogeneous covariance and sample size. Figure 5(a) illustrates the unequal covariance, equal sample size regime ($\rho_{b}>\rho_{r}$, $n_{r}=n_{b}$). Here, we observe that, the segment connecting the population point $\left(\mathcal{R}_{r}(\beta_{1/2}),\mathcal{R}_{b}(\beta_{1/2})\right)$ to the corresponding empirical points $\left(\mathcal{R}_{r}(\tilde{\beta}_{1/2}),\mathcal{R}_{b}(\tilde{\beta}_{1/2})\right)$ and $\left(\mathcal{R}_{r}(\widehat{\beta}_{1/2}),\mathcal{R}_{b}(\widehat{\beta}_{1/2})\right)$ is not parallel to the $45^{\circ}$ line. Instead, we see that it is tilted towards group $b$, reflecting a larger inflation in $\mathcal{R}_{b}$ than $\mathcal{R}_{r}$, on average. This asymmetry is explained by the results of Corollary 2 and Proposition 4, which show that asymmetry in the covariance across groups drives asymmetric contraction of the empirical frontier estimated in finite samples.
In contrast, in the unequal sample size, equal covariance regime ($n_{r}>n_{b}$, $\rho_{r}=\rho_{b}$) illustrated in Figure 5(b), the same segment is approximately parallel to the $45^{\circ}$ line. In other words, with equal covariance, we see that the inflation in $\mathcal{R}_{b}$ than $\mathcal{R}_{r}$ is roughly symmetric for both the known and unknown covariance estimation regimes. Beyond $\lambda=1/2$, we observe that the same pattern holds across the frontier.
Moreover, for both figures, we observe that the two empirical FA frontiers for known and unknown covariances approach each other for $\lambda$ near zero or one, and coincide at the endpoints $\lambda\in\{0,1\}$. The latter is expected, as the endpoints represent the per-group cases in which the optimal estimator does not depend on the knowledge of the covariance. The former observation, that the difference between the two frontiers shrinks around the endpoints, directly reflects our result in Theorem 3, which shows that the bias term of the unknown-covariance estimator includes a factor of $\lambda^{2}(1-\lambda)^{2}$, and therefore vanishes quadratically at the endpoints.